# Current address: ]College of William and Mary, Williamsburg, Virginia 23187, USA Current address: ]Thomas Jefferson National Accelerator Facility, Newport News, Virginia 23606, USA

### v3

@@ I Introduction
The study of nucleon structure has been an active field of research ever since the discovery that nucleons were composite [1, 2, 3]. It is now well known that the nucleon, like all hadrons, is made of partons (quarks, antiquarks, and gluons) and that its structure is dominantly ruled by the strong nuclear interaction described by quantum chromodynamics (QCD) [4, 5]. However, the behavior of the QCD coupling constant [6, 7, 8] determines that our understanding of nucleon structure at long distance is not as developed as that at short distance. At high energy-momentum (short distance) typically well above the GeV scale, , which allows us to treat QCD perturbatively. Presently, the approximations of observables are typically known to order , and some to . On the other hand, at low energy-momentum, , and a perturbative expansion in becomes inapplicable. The problem is further complicated by the rise of complex phenomena, e.g. the confinements of partons, the emergence of hadronic degrees of freedom or the spontaneous breaking of QCD’s near chiral symmetry. In this domain, due to the lack of a mature analytic technique based on the QCD Lagrangian, effective theories or models are often used. In particular, chiral effective field theory (EFT) [9], based on the observed spontaneous breaking of the chiral symmetry in hadronic states, has successfully predicted hadronic observables at low energies.
@@
The principle of EFT is that it describes the nucleon in terms of effective hadronic degrees of freedom instead of the fundamental quark and gluon fields. It is completely non-perturbative in terms of because it employs series whose expansion parameters are based e.g. on the pion mass, the nucleon- mass gap or the characteristic energy scale of the chiral symmetry breaking. However, while leading-order (LO) calculations of EFT are relatively tractable, next-to-leading-order (NLO) and higher order calculations are much more involved. An important experimental task is therefore to provide precise hadron structure data that can test, and potentially guide, EFT calculations and other approaches to non-perturbative QCD. Such data must cover the domain of low momentum transfer squared () where EFT applies, typically below a few tenths of GeV2. Especially important in this context are low nucleon spin structure data [10, 11, 11, 12, 13, 14, 15, 16, 17, 18] which are challenging to obtain experimentally but are crucial for the development of EFT [19]. One such experimental endeavor is that of the Jefferson Lab (JLab) Experimental Group EG4, constituted of experiments E03-006 and E05-111 that used the CEBAF Large Acceptance Spectrometer (CLAS) in Hall B [20]. EG4 was designed to test EFT predictions by precisely measuring the longitudinal spin structure function and its moments for the proton and the deuteuron down to GeV2. These low data, reported in two letters [21, 22], provided benchmarks for EFT calculations. This document is the archival article that presents in detail the EG4 experiments, data analysis procedure for the inclusive scattering channel, results on the proton and the deuteron, and followup analysis and results on the neutron spin structure functions. Additional results on the target- and double-spin asymmetries of pion electroproduction are available [23], but these were focused on the semi-inclusive channel and are not included here.
@@ II Formalism
All formalism presented in this section is based on Ref. [10] or [14], and the algebraic manipulations therein.
@@ II.1 Inclusive Electron Scattering
Unless indicated otherwise, Eq. (3) will be used hereafter for the definition of the invariant mass. Alternatively to , one may characterize the scattering process using the Bjorken scaling variable [24] , or for fixed targets:
@@ II.3 Virtual Photon Cross Sections
The formalism provided in the previous two sections focused on the interaction cross section between the incident electron and the target. It is often useful to also study cross sections for the virtual photon being absorbed by the target, and their dependence on the photon polarization. We follow the formalism of Ref. [5] and Refs. [25, 26].
@@
Using the “relative longitudinal polarization” 111It is called “photon transverse polarization” in Ref. [25]. However since this is the weighting factor of , not , we prefer “relative longitudinal polarization”. of the virtual photon , the total photon absorption cross section can be written as
@@
For the polarized case, one can similarly divide the virtual photon absorption cross section by a spin-1/2 target into several components based on the polarization of the photon, and the total spin of the photon and the target projected along the direction of the photon momentum ( and for the photon spin antiparallel and parallel to the target spin, respectively), and a component resulting from the interference between transverse and longitudinal virtual photon-nucleon amplitudes [5]. From these components one further defines
@@ II.4 Sum Rules and Moments of Structure Functions
In this section we present the Bjorken, GDH, generalized GDH, and generalized polarizability sum rules, which all involve moments of polarized structure functions or, equivalently, and . For reviews of nucleon spin structure and additional sum rules, see Refs. [12, 14, 15, 17].
@@ II.4.1 Bjorken Sum Rule
The Bjorken sum rule, derived for using current algebra and isospin symmetry [27, 28], predates QCD and is not a QCD prediction. However, its generalization to include the -dependence appearing at finite [29, 30, 31] stems from pQCD. Studying the sum rule in the pQCD domain ( GeV2) therefore tests whether QCD correctly describes the strong force when spin degrees-of-freedom are explicit. The Bjorken sum has been actively measured at CERN [32, 33, 34, 35, 36, 37, 38, 39, 40], DESY [41], JLab [42, 43, 44], and SLAC [45, 46, 47, 48, 49, 50], with measurement agreeing to better than 10% with the sum rule expectation [17]. Measurements performed with the higher-energy accelerators, CERN and SLAC, provide data at larger values, offering extensive coverage of the low- part of the Bjorken sum. Measurements at the lower-energy facilities, DESY and JLab, covered the smaller range while overlapping with the SLAC data. Due to their lower energy, DESY and JLab have a limited low- reach, which is supplemented by extrapolating the low- data obtained from CERN and SLAC.
@@
Beside the pQCD domain, the Bjorken sum rule is also important at low due to its close connection to the GDH sum rule (see next section). The behavior of the Bjorken sum at GeV2 was precisely mapped by several experiments at JLab [42, 43, 51, 52]. Note that Eqs. (24-26) are not valid in the low domain and Bjorken sum rule predictions are provided by either models using non-perturbative approaches [53, 54, 55, 56, 57, 58] or EFT calculations [59, 60].
@@ II.4.2 Gerasimov-Drell-Hearn (GDH) Sum Rule
The GDH sum rule [61, 62, 63] is based on dispersion relations, unitarity, relativity, and gauge invariance and is derived for real photoproduction (). Like the Bjorken sum rule, the GDH sum rule is derived in a more general context than that of QCD and also predates it. In particular, the GDH sum rule applies to any type of target. For targets whose internal structure is governed by the strong interaction, such as nucleons and nuclei, the sum rule provides a path to study QCD [64]. For real photons, the GDH sum rule gives
@@
with the pion mass. The anomalous magnetic moment of the target is defined by the total magnetic moment of the particle with the charge of the target in units of the elementary charge, and can be related to the gyromagnetic ratio as . The RHS of Eq. (27) is b for the proton (, and, ), b for the neutron (, , and ), and b for the deuteron (, , and ). Whereas the overall consensus [63] is that the GDH sum rule is theoretically very solid, the question of its validity has been debated, mainly regarding whether decreases fast enough with for the integral to converge. Concurrently, its experimental verification has been the focus of several experiments at MAMI and ELSA [65, 66, 67, 68, 69], BNL [70], JLab [21, 21, 71], GRAAL (ESRF) [72], and HIGS (TUNL) [73] providing results that show, given reasonable assumptions for the large behavior of its integrand, the GDH sum rule is valid for the proton to within the 10% precision of the world data [63, 21, 22, 74].
The GDH integral for the nucleon, when expressed with spin structure functions using Eq. (19), is an integral over [75]. Like the Bjorken sum, the small correction from vanishes when , see Eq. (11). Therefore, the formal expression of the (isovector) GDH and Bjorken integrals are proportional, whereas the domains of applicability of the sum rules are disjoint at high and , respectively. Several methods have been proposed to bridge the two sum rules. Generalization to finite values for the left-hand side (LHS) of Eq. (27) is straightforward since or equivalently and exist also at non-zero . A commonly used definition is [26, 10]:
@@
The above procedure generalizes only the GDH integral but not the GDH sum rule, i.e., the RHS of Eq. (27). Other possible generalizations of the GDH integral as reviewed in Ref. [25] share the same caveat. To generalize the sum rule itself, one needs to note that the integral of can be related to the first forward virtual Compton scattering amplitude [76, 14, 17] using the low energy theorem [77]. Accordingly, a full generalization of the GDH sum rule is [76]:
where the bars and indicate that the elastic contribution (quasi-elastic for nuclear targets) is excluded. In the quark-parton model, this integral gives the quark polarization within the nucleon weighted by their electric charge squared, and provides one of the ingredients for Ellis-Jaffe sum rule [78] and Bjorken sum rule. For the generalized form of the GDH sum rule, Eq. (33), the amplitude can be calculated in EFT if is sufficiently low [79, 80, 81, 82, 83, 59, 84, 85, 60]. The generalized GDH integrals have been measured over an extensive range and on proton, neutron, deuteron and 3He by CERN [86, 32, 33, 34, 35, 36, 37, 87, 88, 38, 39, 40, 89], DESY [90, 91, 41], JLab [92, 93, 94, 95, 96, 51, 97, 98, 99, 100, 101, 102, 103, 21, 22, 71, 104], and SLAC [105, 106, 107, 45, 46, 47, 108, 48, 49, 50]. CERN and SLAC provide the higher coverage, with extensive contributions to the moments from the low- region. JLab data cover the lowest domain. All accelerators provide data points in the intermediate region of a few GeV or less, including the COMPASS experiment [87, 88, 38, 39, 40, 89] for CERN, E143 [105, 106, 107, 45, 46] for SLAC and HERMES for DESY [90, 91, 41].
@@ II.4.3 Forward Spin Polarizability
The interaction of a particle with an electromagnetic field, in our case the reaction of the target particle to the photons exchanged with the electron beam, is dictated at first order in the photon energy by the particle electric charge. If the beam and target are polarized, the target particle’s anomalous magnetic moment also comes into play. Both the electric charge and the magnetic moment reflect the pure elastic reaction expected from a pointlike or perfectly rigid object. At second order, the particle compositeness and deformability must be considered. The deformation of the particle, viz its internal rearrangement under the electromagnetic field, are described by its electromagnetic polarizabilities [109]. The forward spin polarizability arises for the case when the beam and target are polarized longitudinally to the beam direction. Polarizabilities were initially defined for real photons but, just like the GDH sum, they can be generalized to finite , becoming -dependent generalized polarizabilities.
There is no practical method known that allows us to directly measure generalized polarizabilities. Instead, sum rules are used to access them. For , the relevant sum rule involves a higher moment of and [110, 111, 112]:
which can also be calculated in EFT [79, 80, 81, 82, 83, 59, 84, 85, 60]. Equivalently, we can write this integral as:
Measurement of has been carried out for the proton at the real photon point at MAMI [69]. Measurements at small and intermediate have been carried out at JLab on the proton [98, 102, 22] and neutron [19, 98, 103, 104]. A short review of the experimental data and predictions from EFT, MAID [113], and SAID [114] models can be found in Ref. [74].

### v4

@@ I Introduction
The study of nucleon structure has been an active field of research ever since the discovery that nucleons were composite [Frisch:1933, Esterman:1933, Hofstadter:1953zz]. It is now well known that the nucleon, like all hadrons, is made of partons (quarks, antiquarks, and gluons) and that its structure is dominantly ruled by the strong nuclear interaction described by quantum chromodynamics (QCD) [book:QCDhadron, book:Roberts]. However, the behavior of the QCD coupling constant [Prosperi:2006hx, Deur:2016tte, Deur:2023dzc] determines that our understanding of nucleon structure at long distance is not as developed as that at short distance. At high energy-momentum (short distance) typically well above the GeV scale, , which allows us to treat QCD perturbatively. Presently, the approximations of observables are typically known to order , and some to . On the other hand, at low energy-momentum, , and a perturbative expansion in becomes inapplicable. The problem is further complicated by the rise of complex phenomena, e.g. the confinements of partons, the emergence of hadronic degrees of freedom or the spontaneous breaking of QCD’s near chiral symmetry. In this domain, due to the lack of a mature analytic technique based on the QCD Lagrangian, effective theories or models are often used. In particular, chiral effective field theory (EFT) [Bernard:1995dp], based on the observed spontaneous breaking of the chiral symmetry in hadronic states, has successfully predicted hadronic observables at low energies.
@@
The principle of EFT is that it describes the nucleon in terms of effective hadronic degrees of freedom instead of the fundamental quark and gluon fields. It is completely non-perturbative in terms of because it employs series whose expansion parameters are based e.g. on the pion mass, the nucleon- mass gap or the characteristic energy scale of the chiral symmetry breaking. However, while leading-order (LO) calculations of EFT are relatively tractable, next-to-leading-order (NLO) and higher order calculations are much more involved. An important experimental task is therefore to provide precise hadron structure data that can test, and potentially guide, EFT calculations and other approaches to non-perturbative QCD. Such data must cover the domain of low momentum transfer squared () where EFT applies, typically below a few tenths of GeV2. Especially important in this context are low nucleon spin structure data [Anselmino:1994gn, Bass:2004xa, Bass:2004xa, Chen:2005tda, Burkardt:2008jw, Kuhn:2008sy, Chen:2010qc, Aidala:2012mv, Deur:2018roz, Ji:2020ena] which are challenging to obtain experimentally but are crucial for the development of EFT [JeffersonLabE94010:2004ekh]. One such experimental endeavor is that of the Jefferson Lab (JLab) Experimental Group EG4, constituted of experiments E03-006 and E05-111 that used the CEBAF Large Acceptance Spectrometer (CLAS) in Hall B [CLAS:2003umf]. EG4 was designed to test EFT predictions by precisely measuring the longitudinal spin structure function and its moments for the proton and the deuteuron down to GeV2. These low data, reported in two letters [CLAS:2017ozc, CLAS:2021apd], provided benchmarks for EFT calculations. This document is the archival article that presents in detail the EG4 experiments, data analysis procedure for the inclusive scattering channel, results on the proton and the deuteron, and followup analysis and results on the neutron spin structure functions. Additional results on the target- and double-spin asymmetries of pion electroproduction are available [CLAS:2016qxj], but these were focused on the semi-inclusive channel and are not included here.
@@ II Formalism
All formalism presented in this section is based on Ref. [Anselmino:1994gn] or [Kuhn:2008sy], and the algebraic manipulations therein.
@@ II.1 Inclusive Electron Scattering
Unless indicated otherwise, Eq. (3) will be used hereafter for the definition of the invariant mass. Alternatively to , one may characterize the scattering process using the Bjorken scaling variable [Bjorken:1968dy] , or for fixed targets:
@@ II.3 Virtual Photon Cross Sections
The formalism provided in the previous two sections focused on the interaction cross section between the incident electron and the target. It is often useful to also study cross sections for the virtual photon being absorbed by the target, and their dependence on the photon polarization. We follow the formalism of Ref. [book:Roberts] and Refs. [Drechsel:2000ct, Gorchtein:2004jd].
@@
Using the “relative longitudinal polarization” 111It is called “photon transverse polarization” in Ref. [Drechsel:2000ct]. However since this is the weighting factor of , not , we prefer “relative longitudinal polarization”. of the virtual photon , the total photon absorption cross section can be written as
@@
For the polarized case, one can similarly divide the virtual photon absorption cross section by a spin-1/2 target into several components based on the polarization of the photon, and the total spin of the photon and the target projected along the direction of the photon momentum ( and for the photon spin antiparallel and parallel to the target spin, respectively), and a component resulting from the interference between transverse and longitudinal virtual photon-nucleon amplitudes [book:Roberts]. From these components one further defines
@@ II.4 Sum Rules and Moments of Structure Functions
In this section we present the Bjorken, GDH, generalized GDH, and generalized polarizability sum rules, which all involve moments of polarized structure functions or, equivalently, and . For reviews of nucleon spin structure and additional sum rules, see Refs. [Chen:2005tda, Kuhn:2008sy, Chen:2010qc, Deur:2018roz].
@@ II.4.1 Bjorken Sum Rule
The Bjorken sum rule, derived for using current algebra and isospin symmetry [Bjorken:1966jh, Bjorken:1969mm], predates QCD and is not a QCD prediction. However, its generalization to include the -dependence appearing at finite [Kataev:1994gd, Kataev:2005hv, Baikov:2010je] stems from pQCD. Studying the sum rule in the pQCD domain ( GeV2) therefore tests whether QCD correctly describes the strong force when spin degrees-of-freedom are explicit. The Bjorken sum has been actively measured at CERN [SpinMuon:1993gcv, SpinMuonSMC:1997voo, SpinMuonSMC:1994met, SpinMuon:1995svc, SpinMuonSMC:1997ixm, SpinMuonSMC:1997mkb, COMPASS:2010wkz, COMPASS:2015mhb, COMPASS:2016jwv], DESY [HERMES:2002gmr], JLab [Deur:2004ti, Deur:2008ej, Deur:2014vea], and SLAC [E143:1995rkd, E143:1995clm, E154:1997xfa, E143:1998hbs, E154:1997ysl, E155:2000qdr], with measurement agreeing to better than 10% with the sum rule expectation [Deur:2018roz]. Measurements performed with the higher-energy accelerators, CERN and SLAC, provide data at larger values, offering extensive coverage of the low- part of the Bjorken sum. Measurements at the lower-energy facilities, DESY and JLab, covered the smaller range while overlapping with the SLAC data. Due to their lower energy, DESY and JLab have a limited low- reach, which is supplemented by extrapolating the low- data obtained from CERN and SLAC.
@@
Beside the pQCD domain, the Bjorken sum rule is also important at low due to its close connection to the GDH sum rule (see next section). The behavior of the Bjorken sum at GeV2 was precisely mapped by several experiments at JLab [Deur:2004ti, Deur:2008ej, ResonanceSpinStructure:2008ceg, Deur:2021klh]. Note that Eqs. (24-26) are not valid in the low domain and Bjorken sum rule predictions are provided by either models using non-perturbative approaches [Burkert:1992tg, Burkert:1993ya, Soffer:2004ip, Pasechnik:2010fg, Pasechnik:2009yc, Brodsky:2010ur] or EFT calculations [Bernard:2012hb, Alarcon:2020icz].
@@ II.4.2 Gerasimov-Drell-Hearn (GDH) Sum Rule
The GDH sum rule [Gerasimov:1965et, Drell:1966jv, Helbing:2006zp] is based on dispersion relations, unitarity, relativity, and gauge invariance and is derived for real photoproduction (). Like the Bjorken sum rule, the GDH sum rule is derived in a more general context than that of QCD and also predates it. In particular, the GDH sum rule applies to any type of target. For targets whose internal structure is governed by the strong interaction, such as nucleons and nuclei, the sum rule provides a path to study QCD [Deur:2018ntc]. For real photons, the GDH sum rule gives
@@
with the pion mass. The anomalous magnetic moment of the target is defined by the total magnetic moment of the particle with the charge of the target in units of the elementary charge, and can be related to the gyromagnetic ratio as . The RHS of Eq. (27) is b for the proton (, and, ), b for the neutron (, , and ), and b for the deuteron (, , and ). Whereas the overall consensus [Helbing:2006zp] is that the GDH sum rule is theoretically very solid, the question of its validity has been debated, mainly regarding whether decreases fast enough with for the integral to converge. Concurrently, its experimental verification has been the focus of several experiments at MAMI and ELSA [GDH:2001zzk, Ahrens:2006yx, GDH:2003xhc, GDH:2005noz, Dutz:2004zz], BNL [LSC:2008wiu], JLab [CLAS:2017ozc, CLAS:2017ozc, JeffersonLabE97-110:2019fsc], GRAAL (ESRF) [Renard:2000jc], and HIGS (TUNL) [Weller:2002aa] providing results that show, given reasonable assumptions for the large behavior of its integrand, the GDH sum rule is valid for the proton to within the 10% precision of the world data [Helbing:2006zp, CLAS:2017ozc, CLAS:2021apd, Strakovsky:2022tvu].
The GDH integral for the nucleon, when expressed with spin structure functions using Eq. (19), is an integral over [Anselmino:1988hn]. Like the Bjorken sum, the small correction from vanishes when , see Eq. (11). Therefore, the formal expression of the (isovector) GDH and Bjorken integrals are proportional, whereas the domains of applicability of the sum rules are disjoint at high and , respectively. Several methods have been proposed to bridge the two sum rules. Generalization to finite values for the left-hand side (LHS) of Eq. (27) is straightforward since or equivalently and exist also at non-zero . A commonly used definition is [Gorchtein:2004jd, Anselmino:1994gn]:
@@
The above procedure generalizes only the GDH integral but not the GDH sum rule, i.e., the RHS of Eq. (27). Other possible generalizations of the GDH integral as reviewed in Ref. [Drechsel:2000ct] share the same caveat. To generalize the sum rule itself, one needs to note that the integral of can be related to the first forward virtual Compton scattering amplitude [Ji:1999mr, Kuhn:2008sy, Deur:2018roz] using the low energy theorem [Low:1954kd]. Accordingly, a full generalization of the GDH sum rule is [Ji:1999mr]:
where the bars and indicate that the elastic contribution (quasi-elastic for nuclear targets) is excluded. In the quark-parton model, this integral gives the quark polarization within the nucleon weighted by their electric charge squared, and provides one of the ingredients for Ellis-Jaffe sum rule [Ellis:1973kp] and Bjorken sum rule. For the generalized form of the GDH sum rule, Eq. (33), the amplitude can be calculated in EFT if is sufficiently low [Ji:1999pd, Ji:1999sv, Bernard:2002bs, Bernard:2002pw, Kao:2002cp, Bernard:2012hb, Lensky:2014dda, Lensky:2016nui, Alarcon:2020icz]. The generalized GDH integrals have been measured over an extensive range and on proton, neutron, deuteron and 3He by CERN [EuropeanMuon:1987isl, SpinMuon:1993gcv, SpinMuonSMC:1997voo, SpinMuonSMC:1994met, SpinMuon:1995svc, SpinMuonSMC:1997ixm, SpinMuonSMC:1997mkb, COMPASS:2005xxc, COMPASS:2006mhr, COMPASS:2010wkz, COMPASS:2015mhb, COMPASS:2016jwv, COMPASS:2017hef], DESY [HERMES:2000apm, HERMES:1998pau, HERMES:2002gmr], JLab [Amarian:2002ar, JeffersonLabE94-010:2003dvv, CLAS:2002knl, CLAS:2003rjt, RSS:2006tbm, ResonanceSpinStructure:2008ceg, CLAS:2006ozz, CLAS:2008xos, JeffersonLabE01-012:2008kgk, E01-012:2013fky, CLAS:2014qtg, CLAS:2017qga, CLAS:2015otq, CLAS:2017ozc, CLAS:2021apd, JeffersonLabE97-110:2019fsc, E97-110:2021mxm], and SLAC [E142:1996thl, E143:1994vcg, E143:1996vck, E143:1995rkd, E143:1995clm, E154:1997xfa, E155:1999pwm, E143:1998hbs, E154:1997ysl, E155:2000qdr]. CERN and SLAC provide the higher coverage, with extensive contributions to the moments from the low- region. JLab data cover the lowest domain. All accelerators provide data points in the intermediate region of a few GeV or less, including the COMPASS experiment [COMPASS:2005xxc, COMPASS:2006mhr, COMPASS:2010wkz, COMPASS:2015mhb, COMPASS:2016jwv, COMPASS:2017hef] for CERN, E143 [E142:1996thl, E143:1994vcg, E143:1996vck, E143:1995rkd, E143:1995clm] for SLAC and HERMES for DESY [HERMES:2000apm, HERMES:1998pau, HERMES:2002gmr].
@@ II.4.3 Forward Spin Polarizability
The interaction of a particle with an electromagnetic field, in our case the reaction of the target particle to the photons exchanged with the electron beam, is dictated at first order in the photon energy by the particle electric charge. If the beam and target are polarized, the target particle’s anomalous magnetic moment also comes into play. Both the electric charge and the magnetic moment reflect the pure elastic reaction expected from a pointlike or perfectly rigid object. At second order, the particle compositeness and deformability must be considered. The deformation of the particle, viz its internal rearrangement under the electromagnetic field, are described by its electromagnetic polarizabilities [Hagelstein:2015egb]. The forward spin polarizability arises for the case when the beam and target are polarized longitudinally to the beam direction. Polarizabilities were initially defined for real photons but, just like the GDH sum, they can be generalized to finite , becoming -dependent generalized polarizabilities.
There is no practical method known that allows us to directly measure generalized polarizabilities. Instead, sum rules are used to access them. For , the relevant sum rule involves a higher moment of and [Gell-Mann:1954ttj, Guichon:1995pu, Drechsel:2002ar]:
which can also be calculated in EFT [Ji:1999pd, Ji:1999sv, Bernard:2002bs, Bernard:2002pw, Kao:2002cp, Bernard:2012hb, Lensky:2014dda, Lensky:2016nui, Alarcon:2020icz]. Equivalently, we can write this integral as:
Measurement of has been carried out for the proton at the real photon point at MAMI [Dutz:2004zz]. Measurements at small and intermediate have been carried out at JLab on the proton [CLAS:2008xos, CLAS:2017qga, CLAS:2021apd] and neutron [JeffersonLabE94010:2004ekh, CLAS:2008xos, CLAS:2015otq, E97-110:2021mxm]. A short review of the experimental data and predictions from EFT, MAID [Drechsel:1998hk], and SAID [Workman:2012jf] models can be found in Ref. [Strakovsky:2022tvu].